Displaying report 1-1 of 1.
Reports until 09:45, Thursday 05 December 2024
Equipment General
Torrey Cullen - posted 09:45, Thursday 05 December 2024 - last comment - 17:16, Friday 06 December 2024(12034)
Power Meter Repair

[Ian, Torrey]

While working in the lab, the ND filter attached to the power meter would not slide into place. Upon inspection we noticed that the sensor was slightly askew . We took it apart and found that the glue used to glue the sensor to the plastic back had come apart (they used a miniscule amount for some reason). We glued the sensor back to the plastic casing and cleaned the ND filter. Put things back together and things are working has intended.

Additionally we went to look at the second power meter Alex uses and the ND slider was not working on that one either. We will repair this one as well, have not gotten to it yet.

Images attached to this report
Comments related to this report
Torrey Cullen - 17:16, Friday 06 December 2024 (12039)

Took apart and re-glued the one alex uses. It is drying under some alluminum foil in B111B for the time being. Will reassemble Monday.

Displaying report 1-1 of 1.